OpenStudy (anonymous):

What is and how do you simplify (3/x) - (3/x+1) +(3/x+2)? any restrictions?

OpenStudy (anonymous):

\[\frac{ 3 }{ x } - \frac{ 3 }{ x+1 } + \frac{ 3 }{ x+2 }\]

OpenStudy (primeralph):